OpenStudy (anonymous):

405u^10x^6/75u^3x^4 Simplify by removing factors of 1

OpenStudy (anonymous):

\[\frac{405u^{10}x^6}{75u^3x^4}\] subtract the exponents

OpenStudy (anonymous):

if anyone can help I would appreciate it. I am soooo confused. IF I could be walked through it I would appreciate it too.

OpenStudy (anonymous):

get \[\frac{405}{75}6^{10-3}x^{6-4}\] \[5.4u^7x^2\]

OpenStudy (anonymous):

i made a typo there hold on

OpenStudy (anonymous):

first of all just divide \[\frac{405}{75}=\frac{27}{5}=5.4\]

OpenStudy (anonymous):

then subtract the exponents. you have \[\frac{u^{10}}{u^3}=u^{10-3}=u^7\]

OpenStudy (anonymous):

and also \[\frac{x^6}{x^4}=x^{6-4}=x^2\]

OpenStudy (anonymous):

so answer is \[\frac{27}{5}u^7x^2\]

OpenStudy (anonymous):

all those parenthesis confuse me. lol give me a second to look that over please.

OpenStudy (anonymous):

take your time. but idea is to subtract the exponents of the like terms

OpenStudy (anonymous):

can you write that answer not from the editor? I would like to see how it is supposed to be written and all the \ \ confuse me. It does make sense to me when answers look like this : 27(5)u^7x^2 is that how it is supposed to look?

OpenStudy (anonymous):

if you are not seeing the text properly, refresh your browser. i bet it will work

OpenStudy (anonymous):

you should see just the nice expressions, not the editor. try it and let me know

OpenStudy (anonymous):

you should not see any \\ or ] or }

OpenStudy (anonymous):

i see a whole lot of those. just refreshed. the same. hold on let me get out of explorer.

OpenStudy (anonymous):

there we go.

OpenStudy (anonymous):

thank you... I think I may have it but if I get a similar one I may post that as well just to walk through it again.